### v4

@@ 8.3.4 The Lack of Proactive Defenses
Most existing defense approaches, such as safety alignment and adversarial training, are passive in nature, focusing on fortifying models against potential attacks. However, proactive defenses, which actively counter potential attacks before they succeed, have received limited attention in the literature. For example, a proactive defense against model extraction could involve poisoning or injecting backdoors into extraction attempts, rendering the extracted model unreliable. Another approach could be to provide deliberately nonsensical or easily detectable responses when a malicious user requests advice for illegal activities, such as planning a robbery. Such proactive strategies could serve as powerful deterrents to potential attackers. However, designing effective proactive defenses for different types of safety threats remains an open challenge and a promising direction for future research.
8.3.5 Detection Has Been Overlooked in Current Defenses
Detection methods play a crucial role in identifying potential vulnerabilities and abnormal behaviors in models, effectively acting as active monitors. When integrated with other defense mechanisms, detection systems can trigger automatic safety responses whenever a model behaves unexpectedly or generates harmful outputs. Despite their importance, however, existing defense strategies have largely overlooked the integration of detection systems within their pipelines. By combining detection with other safety measures, it becomes possible to develop more resilient models capable of dynamically responding to emerging threats. For example, stronger attacks may be more easily detected, providing an opportunity for proactive defense. An open question remains: What is the most effective way to integrate detection as a core component of a defense system, and how can detection and other defense mechanisms complement and enhance each other?
8.3.6 The Current Data Usage Practices Must Change
The current data usage practices in the AI development lifecycle are neither sustainable nor ethical. We identify three major issues with how data is currently used:
  1. 1
    Lack of Consent and Recognition: Many large models are trained on web-crawled data without the consent of data owners or formal recognition or reward for their contributions. This practice raises significant ethical and legal issues.
  2. 2
    The Explosion of Generated Data: A large volume of generated data has been uploaded to the internet, much of which is fake, toxic, or otherwise harmful. However, there is no clear system in place to identify which model created this data or who was responsible for its generation.
  3. 3
    Depletion of “Free” Data: The “free” data available on the internet is rapidly diminishing. Users are becoming less motivated or outright refusing to contribute valuable data, especially when their contributions are neither acknowledged nor rewarded.
To address these challenges, the AI industry must establish a healthy and sustainable data ecosystem where data contributors are recognized and rewarded. Achieving this requires answering the following key questions:
  • Who Used My Data? This question addresses the protection of copyright for original training data, commonly known as membership inference. Membership inference techniques are essential to determine whether specific samples were included in a model’s training data. Such methods would empower data owners to verify how their data is being used and protect their legitimate rights and interests.
  • Who Generated the Data? This question focuses on protecting the copyright of generated data, a process referred to as model attribution. Techniques for model attribution are needed to identify the model responsible for generating a given sample (such as text, image, video, or audio). Model attribution should also include identifying the user responsible for creating the content, including metadata such as IDs or usernames. These technologies would not only protect the copyright of generated content but also encourage responsible data generation by ensuring that harmful or malicious content can be traced back to its creator.
  • Which Samples Contribute to a Generated Output? This question pertains to data attribution in the context of AIGC. Every piece of generated content can be seen as a combination of “inspirations” drawn from a specific set of training samples. Data attribution techniques are needed to identify the training samples that most significantly contributed to the generated content. Contributors whose samples play a critical role in shaping the content should receive a share of the profits if the content is used commercially, promoting a fair and transparent data economy.
Beyond these questions, many other critical issues must be addressed to create a sustainable and ethical data ecosystem. Ensuring that data contributors are fairly acknowledged and incentivized is essential for promoting accountability and transparency in AIGC.
@@ 8.3.7 Safe Embodied Agents
Most safety threats studied today are digital. However, as embodied AI agents are increasingly deployed in the physical world, new physical threats will emerge, potentially causing real harm and loss to humans. Ensuring the safety of these embodied agents has therefore become a critical concern. Safe agents must be resilient to adversarial inputs, capable of self-regulating harmful behaviors, and consistently aligned with human values.
Achieving this requires deeply integrating safety mechanisms into the agents’ decision-making processes, enabling them to handle unexpected challenges while maintaining robustness and reliability. The primary challenge is designing safety protocols that allow these agents to perform complex tasks autonomously, while ensuring they remain trustworthy and safe in dynamic and unpredictable environments. As agents gain greater autonomy, ensuring their safety becomes not only a technical challenge but also a significant ethical responsibility.
@@ 8.3.8 Safe Superintelligence
As AI progresses toward AGI and superintelligence, embedding inherent safety mechanisms into large models to ensure predictable, value-aligned behavior becomes a critical challenge. While the technical path to safe superintelligence remains uncertain, several promising mechanisms offer potential solutions:
  • Oversight System: One system cannot be both superintelligent and trustworthy, like humans. However, we can develop an oversight system to monitor and regulate the primary system’s behavior, intervening when necessary. A key challenge is to ensure the reliability and robustness of the oversight system itself. This leads to the Oversight Paradox: If a superintelligent AI is monitored by another AI (the oversight system), the oversight system must be at least as capable as the superintelligent AI to reliably detect and prevent undesirable behaviors. However, this raises the question: who monitors the oversight system to ensure it doesn’t fail or act contrary to its purpose?
  • Safety Layer: This approach embeds a dedicated safety layer [520, 521] directly within the model’s architecture, acting as a gatekeeper to filter outputs against predefined safety constraints. Such layers could be dynamically updated based on real-time feedback or optimized for specific tasks.
  • Safety Expert: This approach incorporates specialized “safety experts" within the Mixture of Experts (MoE) framework [522, 523, 524, 525] to handle safety-critical tasks. By dynamically routing high-stakes queries to these experts, safety considerations are prioritized in decision-making. The recurring challenge is developing a truly safe expert model that can consistently perform as expected in all scenarios.
  • Adversarial Alignment: This approach leverages adversarial safety principles to align models with human values. It involves training models to exploit vulnerabilities in existing safety mechanisms, then refining these mechanisms to resist adversarial prompts. Despite its promise, challenges such as high computational costs and the risk of unintended behaviors remain significant concerns.
  • Safety Consciousness: This approach involves embedding a safety-conscious framework into the model’s foundational training to promote ethical reasoning and value alignment as intrinsic behaviors. The goal is to make safety a core characteristic, enabling the model to dynamically adapt to diverse and evolving scenarios. Safety consciousness may be formulated as a type of safety tendency: an inherent inclination to generate low-risk responses and shape outputs based on an awareness of potential harmful consequences, similar to human decision-making processes.
@@ 8.4 A Call for Collective Action
Safeguarding large models against adversarial manipulation, misuse, and harm is a global challenge that requires the collective efforts of researchers, practitioners, and policymakers. The following sections outline a research agenda aimed at advancing large-model safety through collaboration and innovation.
@@ 8.4.1 Defense-Oriented Research
Current research on large-model safety is heavily skewed toward attack strategies, with significantly fewer efforts dedicated to developing defense mechanisms. This imbalance is concerning, as the sophistication of attacks continues to outpace the development of effective defenses. To address this gap, we advocate for a shift in research priorities toward defense strategies. Researchers should not only investigate attack mechanisms but also focus on developing robust defenses to mitigate or prevent these threats. A balanced approach is crucial for advancing the field of safety research.
Moreover, future defense research should emphasize integrated approaches. New defense methods should not be proposed or implemented in isolation but rather integrated with existing approaches to build cumulative protection. Defense research should be viewed as a continuous, evolving effort, where new methods are layered onto established ones to enhance overall effectiveness. However, the diversity of defense strategies presents a challenge. Developing frameworks to effectively incorporate different defense mechanisms will require a collective effort from the research community.
@@ 8.4.2 Dedicated Safety APIs
To support research and testing, commercial AI models should offer a dedicated safety API. This API would allow researchers to assess and enhance the safety of these models by subjecting them to a variety of adversarial and safety-critical scenarios. By providing such an API, commercial providers can enable external safety evaluations without disrupting the general services offered to users. This would foster collaboration between industry and academia, facilitating continuous improvement in model safety.
@@ 8.4.3 Open-Source Platforms
The AI safety community would greatly benefit from the development and open-source release of safety platforms and libraries. These tools would facilitate the rapid evaluation, testing, and improvement of safety mechanisms across a variety of models and applications. Open-sourcing these platforms would foster collaboration and transparency, enabling researchers and practitioners to share best practices, benchmark safety techniques, and contribute to the establishment of universal safety standards.
@@ 8.4.4 Global Collaborations
The pursuit of AI safety is a global challenge that transcends national borders, requiring coordinated efforts from academia, technology companies, government agencies, and non-profit organizations. Effective collaboration on a global scale is essential to addressing the potential risks associated with advanced AI systems. By fostering international cooperation, we can more efficiently tackle complex safety issues and establish unified standards that guide the safe development and deployment of AI technologies.
@@
  • International Safety Alliances: Establishing global alliances dedicated to AI safety can bring together experts and resources from around the world. These alliances would focus on sharing research findings, coordinating safety evaluations, and developing universal safety benchmarks that reflect diverse regional needs and values.
  • Cross-Border Data Sharing Frameworks: Access to diverse datasets is essential to improving the robustness and fairness of AI models. Developing secure and ethical frameworks for cross-border data sharing would allow researchers to test models across a wide range of scenarios and ensure that safety mechanisms are universally applicable.
  • Joint Research Programs: Collaborative research programs that unite academic institutions, industry leaders, and government agencies can drive innovation in AI safety. These programs should focus on areas such as adversarial defense, safety alignment, and real-time adaptability, ensuring that their findings are broadly applicable across various AI systems.
  • Global Safety Competitions and Challenges: Building on the concept of open safety competitions, international challenges could be organized to engage the brightest minds from around the world. These competitions would address critical safety issues, encourage the development of innovative solutions, and foster a sense of shared responsibility in advancing AI safety.
  • Policy and Regulatory Harmonization: A collaborative approach to AI governance can help align safety regulations and policies across countries. This harmonization would prevent the misuse of AI technologies while promoting responsible development and deployment practices globally.
Global collaborations not only enhance the effectiveness of AI safety research but also promote transparency, trust, and accountability in the development of advanced AI systems. By working together across borders and disciplines, we can ensure that AI technologies benefit humanity while minimizing the risks associated with their deployment.
@@ 9 Conclusion
In this paper, we surveyed 390 technical papers on large model safety, encompassing Vision Foundation Models (VFMs), Large Language Models (LLMs), Vision-Language Pretraining (VLP) models, Vision-Language Models (VLMs), Diffusion Models (DMs), and large-model-powered agents. We provided a comprehensive taxonomy of threats and defenses, highlighting the evolving challenges these models face. Despite notable progress, numerous open challenges remain, particularly in understanding the fundamental vulnerabilities of large models, establishing comprehensive safety evaluations, developing scalable and effective defense mechanisms, and ensuring sustainable data practices. More importantly, achieving safe AI will require collective efforts from the global research community and international collaboration. We hope this paper could serve as a useful resource for researchers and practitioners, driving ongoing efforts to build safe, robust and trustworthy large-scale models.

### v5

@@ 8.3.3 The Lack of Proactive Defenses
Most current defense approaches, such as safety alignment and adversarial training, are passive, aiming to safeguard models against incoming attacks. In contrast, proactive defenses, which anticipate and counter attacks before they succeed, remain underexplored. For instance, proactively defending against model extraction might involve poisoning or backdooring extraction attempts to make the stolen model unusable, or providing deliberately nonsensical or easily flagged responses when users seek illegal advice. Such proactive strategies could be powerful deterrents. However, designing effective proactive defenses for diverse safety threats remains an open challenge and an important avenue for future research.
8.3.4 Detection Is Overlooked in Current Defenses
Detection methods are essential for identifying potential vulnerabilities and abnormal behaviors in models, serving as active monitors. When combined with other defense mechanisms, detection systems can automatically trigger safety interventions whenever a model behaves unexpectedly or generates harmful outputs. Despite their importance, most current defense strategies have not fully integrated detection into their pipelines. Recent proposal such as chain-of-thought (CoT) monitoring offer even deeper insight into model reasoning by tracking the intermediate steps and thought processes leading to a model’s decision [700]. By monitoring CoT outputs, it becomes possible to detect early signs of unsafe or manipulative reasoning, enabling more timely and targeted safety interventions.
Integrating detection with other safety measures enables the development of more robust models that can dynamically respond to emerging threats. For instance, stronger or novel attacks may be more readily detected, allowing for timely, proactive defenses. An open question remains: What is the most effective way to make detection, including chain-of-thought monitoring, a core component of defense systems, and how can detection and other defense mechanisms best complement and enhance each other?
@@ 8.3.5 Safe Embodied Agents
Most safety threats studied today are primarily digital. However, as embodied AI agents are increasingly deployed in the physical world, new forms of physical threats will emerge—threats that can result in tangible harm or loss to humans. Ensuring the safety of embodied agents has therefore become a critical priority. Safe agents must demonstrate resilience to adversarial inputs, possess mechanisms for self-regulation against harmful behaviors, and maintain consistent alignment with human values.
Achieving this requires deeply embedding safety mechanisms into agents’ decision-making processes at every step, enabling them to handle unexpected challenges while maintaining robustness and reliability. The key challenge lies in designing safety protocols that empower agents to perform complex tasks autonomously, while remaining trustworthy and safe in dynamic, unpredictable environments. As agents gain greater autonomy, ensuring their safety becomes not only a technical hurdle but also a significant ethical responsibility.
@@ 8.3.6 Safe Superintelligence
As AI advances toward AGI and superintelligence, embedding intrinsic safety mechanisms into large models to ensure predictable, value-aligned behavior is a critical challenge. Although the technical roadmap for achieving safe superintelligence (SSI) remains uncertain, several promising approaches offer potential solutions:
  • Oversight System: No single system, human or AI, can be both superintelligent and inherently trustworthy. To address this, an external oversight system can be designed to monitor and regulate the primary system’s behavior, intervening when necessary. The main challenge lies in ensuring the oversight system’s own reliability and trustworthiness. This gives rise to the Oversight Paradox: The oversight system must be at least as intelligent as, or even more intelligent than, the system it oversees; otherwise, it risks being easily deceived by the system it is meant to monitor. This, in turn, prompts the question: Who monitors the oversight system to guarantee it doesn’t act against its intended purpose?
  • Safety Switch: The safety switch is an emergent “stop button” mechanism designed to immediately shift a model into an ultra-safe operational mode when necessary. One implementation is the integration of a dedicated safety layer [701, 702] directly into the model’s architecture. Beyond safety layers, a safety switch could be realized through runtime overrides, policy re-routing, or external supervisory triggers, all enabling rapid response to unforeseen risks. The key goal is to provide a robust, flexible mechanism that can be activated to prioritize safety above all else, adapting dynamically to real-time feedback and evolving contexts.
  • Safety Expert Model: This strategy introduces specialized safety expert models into the Mixture of Experts (MoE) framework [703, 704, 705, 706] to manage safety-critical tasks. By dynamically routing high-risk or sensitive queries to these experts, the system ensures that safety considerations are prioritized in decision-making. The main challenge remains developing expert models that can consistently and reliably uphold safety across diverse scenarios.
  • Adversarial Alignment: This approach employs adversarial safety principles to better align models with human values. It trains models to identify and exploit weaknesses in current safety mechanisms, which are then iteratively improved to withstand adversarial prompts. While this method shows promise, it also faces notable challenges, including high computational demands and the risk of introducing unintended behaviors.
  • Safety Consciousness: This approach embeds a safety-aware framework into the model’s foundational training, fostering ethical reasoning and value alignment as intrinsic behaviors. The aim is to make safety a core characteristic, enabling the model to dynamically adapt to diverse and evolving scenarios. Safety consciousness can be viewed as a form of safety tendency: an inherent inclination to produce low-risk responses and shape outputs with an awareness of potential harm, much like human decision-making.
@@ 8.4 A Call for Collective Action
Safeguarding large models from adversarial manipulation, misuse, and harm is a global challenge that demands coordinated efforts from researchers, practitioners, and policymakers. The following sections present a research agenda designed to advance large-model safety through collaboration and innovation.
@@ 8.4.1 Defense-Oriented Research
Current research on large-model safety is heavily skewed toward attack strategies, with far less emphasis on developing defenses. This imbalance is concerning as attack sophistication continues to outpace effective safeguards. To address this, we advocate for a shift in research priorities toward robust defense development. Researchers should focus not only on attack mechanisms, but also on practical and preventative defenses to mitigate emerging threats. A balanced approach is essential for advancing safety.
Future defense research should also emphasize integration. New methods should be combined with existing approaches to build layered, cumulative protection as defense is a continuous, evolving process. However, the diversity of defense strategies makes integration challenging, underscoring the need for community-driven frameworks capable of effectively combining multiple defense mechanisms—a truly comprehensive “super safety” framework.
@@ 8.4.2 Dedicated Safety APIs
To facilitate research and testing, commercial AI models should provide a dedicated safety API. Such an API would enable researchers to evaluate and strengthen model safety by exposing models to diverse adversarial and safety-critical scenarios. By making this functionality available, commercial providers can support external safety assessments without impacting regular user services. This approach would foster industry-academia collaboration and drive ongoing improvements in model safety.
@@ 8.4.3 Open-Source Platforms
The AI safety community would benefit greatly from the development and open-source release of safety platforms and libraries. Such tools would accelerate the evaluation, testing, and enhancement of safety mechanisms across diverse models and applications. Open-sourcing these resources would promote collaboration and transparency, allowing researchers and practitioners to share best practices, benchmark safety solutions, and contribute to the creation of universal safety standards.
@@ 8.4.4 Global Collaborations
The pursuit of AI safety is a global challenge that transcends national borders, requiring coordinated efforts from academia, industry, government agencies, and non-profit organizations. Effective international collaboration is essential to address the risks posed by advanced AI systems. By fostering global cooperation, we can more effectively tackle complex safety challenges and establish unified standards to guide the responsible development and deployment of AI technologies.
@@
  • International Safety Alliances: Forming global alliances dedicated to AI safety can unite experts and resources worldwide. These alliances would facilitate sharing research insights, coordinating safety assessments, and developing universal safety benchmarks that account for diverse regional needs and values.
  • Cross-Border Data Sharing: Access to diverse datasets is crucial for enhancing the robustness and fairness of AI models. Establishing secure and ethical frameworks for cross-border data sharing would enable researchers to evaluate models in a broader range of scenarios, ensuring that safety mechanisms are effective and universally applicable.
  • Joint Safety Research Programs: Collaborative research initiatives uniting academic institutions, industry leaders, and government agencies can drive innovation in AI safety. These programs should prioritize areas such as risk prediction, the development of safety guardrails, model enhancement, and safe reasoning strategies, ensuring that their findings are broadly applicable to a wide range of AI systems.
  • International Safety Competitions: Expanding on the concept of open safety competitions, international challenges can be organized to engage top talent worldwide. These competitions would tackle critical and long-term safety issues, drive the development of innovative solutions, and promote a shared sense of responsibility for advancing AI safety.
  • Policy and Regulatory Implementation: Effective AI governance requires practical, enforceable mechanisms. To this end, we advocate for the development of specialized agent systems capable of automatically auditing AI models for compliance with relevant regulations and policies. Bridging the gap between policy and technology is essential for advancing trustworthy and responsible AI, ensuring that high-level regulations can be systematically applied and verified in real-world deployments.
Global collaboration not only strengthens the effectiveness of AI safety research but also promotes transparency, trust, and accountability in the development of advanced AI systems. By working together across borders and disciplines, we can ensure that AI technologies deliver broad benefits to humanity while minimizing associated risks.
@@ 9 Conclusion
In this paper, we conducted a comprehensive survey of safety research on Vision Foundation Models (VFMs), Large Language Models (LLMs), Vision-Language Pretraining (VLP) models, Vision-Language Models (VLMs), Diffusion Models (DMs), and large model powered Agents. We presented a comprehensive taxonomy of existing threats and defenses, highlighting the evolving challenges these models face. Despite significant progress, many open challenges remain, particularly in understanding the fundamental vulnerabilities of large models, establishing robust safety evaluation protocols, and developing scalable, proactive, and integrated defense. Achieving safe AI will require not only technical advances but also collective action from the global research community and international collaboration. We hope this paper serves as a valuable resource for researchers and practitioners, helping to drive ongoing efforts to build safe, robust, and trustworthy large-scale AI systems.
10 Author Contributions
Xingjun Ma designed the survey structure, organized the review process, wrote the challenges and conclusion sections, and prepared the final manuscript. All authors discussed the outline, contributed to drafting and revision, and approved the final manuscript. Yu-Gang Jiang initiated the project, secured resources, guided scientific direction, coordinated teams, and supervised final review and submission.
Vision Foundation Model Safety  Ye Sun and Hanxun Huang surveyed visual backbones, scalable pre-training strategies, and key safety studies, and wrote the draft of this section. James Bailey, Jingfeng Zhang, Yiming Li, Mingming Gong, Tongliang Liu, Shirui Pan, and Sarah Erfani provided expertise in adversarial robustness, efficient architecture design, and graph signal processing, and reviewed this section.
Large Language Model Safety   Yixu Wang, Yifan Ding, Yige Li, Haonan Li, Xudong Han, and Xiang Zheng covered alignment, jailbreaks, prompt injection, and extraction threats, and prepared the draft of this section. Xipeng Qiu, Tim Baldwin, Xiangyu Zhang, Neil Gong, and Yang Liu advised on multilingual scaling, generalization theory, privacy, and alignment, and refined the manuscript.
Vision-Language Pre-training Model Safety  Xin Wang and Jiaming Zhang reviewed literature on large-scale corpora, pre-training objectives, and transfer evaluation, and prepared the initial draft of this section. Tianwei Zhang, Jindong Gu, and Siheng Chen provided guidance on secure deployment, domain generalization, and representation learning, and further refined the section.
Vision-Language Model Safety  Ruofan Wang and Zuxuan Wu reviewed cross-modal architectures, datasets, and open challenges, and prepared the initial draft of this section. Dacheng Tao, Shiqing Ma, Cong Wang, Yang Zhang, and Masashi Sugiyama contributed expertise in multimodal defense and safety alignment and provided feedback on benchmarks and deployment.
Diffusion Model Safety  Yifeng Gao designed the structure of this section and reviewed literature on adversarial attacks, jailbreaks, backdoors, and intellectual property protection. Hengyuan Xu, Yunhan Zhao, and Yunhao Chen surveyed research on membership inference and data/model extraction attacks. Chaowei Xiao, Baoyuan Wu, Tianyu Pang, Yinpeng Dong, and Cihang Xie provided technical guidance on generative model robustness and critically revised this section.
Agent Safety  Yutao Wu designed the structure of this section and prepared the initial draft. Bo Li, Yang Zhang, Liu, Ruoxi Jia, Cong Wang, Yang Liu, Siheng Chen, and Chaowei Xiao contributed insights on indirect prompt injection attacks and defenses, secure tool use, and helped refine the section’s structure.